When Mr. Tan was his son's present age, his son was only 4 years old. When his son reaches his present age, Mr. Tan will be 79 years old. How old is each of them now?
Answer by Edwin McCravy(20055)

When Mr. Tan was his son's present age, his son was only 4 years old.  When his
son reaches his present age, Mr. Tan will be 79 years old. How old is each of
them now?

Let Mr. Tan's age be T.  
Let his son's age be S.

When Mr. Tan was his son's present age, his son was only 4 years old.

Let y = the number of years ago it was when Mr. Tan was his son's 
present age. So we have two equations:

T - y = S, and S - y = 4

When his son reaches his present age, Mr. Tan will be 79 years old. 

Let z = the number of years in the future it will be when his son
reaches Mr. Tan's age.  So we have two more equations:

S + z = T, and T + z = 79.

We have four equations in four unknowns.

(1)  T - y = S
(2)  S - y = 4
(3)  S + z = T
(4)  T + z = 79

Eliminate y from (1) and (2), and eliminate z from (3) and (4).
Then solve the resulting two equations for T and S.

Answer:

S = 29,   T = 54,   y = 25,   z = 25.

Perhaps you could have seen right away that y and z are the same.
Then you could have used only 3 unknowns.  And perhaps you could
have seen that y and z were just the difference in their ages. If 
so, then you could have used only 1 equation and 1 unknown.  But 
it's easier to set up by using 4 equations and 4 unknowns.

Edwin